OpenStudy (ja1):

Check: Solve 3x2 + 4x + 2 = 0. Round solutions to the nearest hundredth. x ≈ −0.59 and x ≈ −3.41 x ≈ 0.20 and x ≈ 1.14 x ≈ −0.20 and x ≈ −1.14 No real solutions

OpenStudy (ja1):

\[x = \frac{ -4 \pm \sqrt{4^{2} - 4(3)(2)} }{ 2(3) }\]

OpenStudy (raden):

check the discriminant value (D), if D < 0 then no real solutions with D = b^2 - 4ac

OpenStudy (luigi0210):

Use the discriminant to determine is you can solve it: \[\sqrt{4^2-4(3)(2)}\] \[\sqrt{16-24}\]

OpenStudy (ja1):

\[x = \frac{ -4 \pm \sqrt{16 - 24} }{ 2(3) }\]

OpenStudy (ja1):

\[x = \frac{ -4 \pm \sqrt{-8} }{ 2(3) }\]

OpenStudy (ja1):

\[x = \frac{ -4 \pm 2.82 }{ 6 }\]

OpenStudy (ja1):

Right so far?

OpenStudy (ja1):

\[x = \frac{ -4 + 2.82 }{ 6 }\] \[x = \frac{ -4 - 2.82 }{ 6 }\]

OpenStudy (raden):

sqrt(-8), it is a complex number

OpenStudy (ja1):

Yes but you would shorten it and do \[\approx\] right?

OpenStudy (luigi0210):

When you have a negative in a radical you get imaginary solutions.

OpenStudy (ja1):

\[x \approx \frac{ -1.18 }{ 6 }\] \[x \approx \frac{ -6.82 }{ 6 }\]

OpenStudy (ja1):

\[x \approx 0.196\] \[x \approx -1.136\]

OpenStudy (ja1):

So this would be "No real solution" right?

OpenStudy (ja1):

I am confused <_<

OpenStudy (luigi0210):

Yes, there are no real solutions.. and oh hey JA didn't realize it was you! :D

OpenStudy (ja1):

lol hey :)

OpenStudy (ja1):

Wait but how could you know before solving it?

OpenStudy (luigi0210):

By checking the discriminant

OpenStudy (ja1):

? I am very bad with mathematical terms :/

OpenStudy (luigi0210):

The \[\sqrt{b^2-4ac}\] :P

OpenStudy (ja1):

Yes yes but how would you see which part tells you?

OpenStudy (luigi0210):

If it's positive then it'll have real solutions If it's negative it has imaginary solutions

OpenStudy (ja1):

OHHHHHH I see now

OpenStudy (ja1):

Blind as a bat <_<

OpenStudy (luigi0210):

Naw :P

OpenStudy (ja1):

lol k thanks a heap :)

OpenStudy (luigi0210):

Yup!